Find orthogonal trajectory to the curve given by π‘Ÿ = π‘Ž(1 + cos πœƒ)

1
Expert's answer
2021-12-30T10:42:21-0500
"r = a(1 + \cosθ)"

Differentiate


"dr=-a\sin \theta d\theta"

"a=\dfrac{r}{1+\cos \theta}"

"dr=-\dfrac{r}{1+\cos \theta}\sin \theta d\theta"

"-\dfrac{1}{r}\dfrac{dr}{d\theta}=\dfrac{\sin \theta}{1+\cos \theta}"

This is the differential equation of the direction field "D_1" for the given family "F_1." To find the differential equation of the direction field orthological to "D_1," we replace "\dfrac{dr}{d\theta}" by "-r^2\dfrac{d\theta}{dr}."

The differential equation for the orthogonal trajectories becomes


"r\dfrac{d\theta}{dr}=\dfrac{\sin \theta}{1+\cos \theta}=\dfrac{2\sin (\dfrac{\theta}{2})\cos (\dfrac{\theta}{2})}{2\cos^2 (\dfrac{\theta}{2})}=\tan(\dfrac{\theta}{2})"

which is a case of variable-separable, and on integration gives


"\ln r=2\ln(\sin(\dfrac{\theta}{2}))+\ln(2c)"

This is the equation of the orthogonal family "F_2."

Since "r = c(1 – \cosθ)" represents the same curve as "r = c(1 + \cosθ)," the member of "F_2" with label "c" is the same as the member of "F_1" with label "a = \u2013c." Thus the given family is selforthogonal.